OpenStudy (jasontodd):

What is the x-intercept of the line with this equation 1/3x+y=−15 ? Enter your answer in the box. (__,0)

OpenStudy (eliesaab):

Solve for x, when y=0

OpenStudy (jasontodd):

2?

OpenStudy (solomonzelman):

` ` Example. Find the x-intercept of the line with the equation, \(\color{black}{\displaystyle (3/5)y+(4/5)x=1 }\). ` ` Solution. You are looking for an x-intercept, so set \(\color{}{y=0}\), and solve for x. \(\color{black}{\displaystyle (3/5){\color{red}{\times 0}}+(4/5)x=1 }\) \(\color{black}{\displaystyle (4/5)x=1 }\) \(\color{black}{\displaystyle (4/5)x{\color{red}{\times (5/4)}}=1{\color{red}{\times (5/4)}} }\) \(\color{black}{\displaystyle x=5/4}\). So, your x-intercept is at x=5/4. (Or, if you want to write a coordinate/point, then \(\color{blue}{(5/4,~~0)}\).)